OpenStudy (anonymous):

substitution method -7x+2y=18 6x+6y=0

hero (hero):

6x + 6y = 0 3(2x + 2y) = 0 2x + 2y = 0 7x - 2y = 0 9x = 0 x = 0 -7(0)+2y = 18 2y = 18 y = 18/2 y = 9